A phone company offers two monthly plans. Plan A costs
$21
plus an additional
$0.14
for each minute of calls. Plan B costs
$8
plus an additional
$0.18
for each minute of calls.
For what amount of calling do the two plans cost the same?
What is the cost when the two plans cost the same?

Answer by Theo(13342)

plan A equation is C = 21 + .14 * x

plan B equation is C = 8 + .18 * x

C is the cost of each plan.

when the cost of each plan is the same, then 21 + .14 * x = 8 + .18 * x

subtract 8 from both sides of the equation subtract .14 from both sides of the equation to get:

21 - 8 = .18 * x - .14 * x

simplify to get 13 = .04 * x

divide both sides of the equation by .04 to get:

x = 13 / .04 = 325

the plans will cost the same when the number of minutes used each month is equal to 325.

when that happens:

plan A costs 21 + .14 * 325 = 66.5

plan B cost 8 + .18 * 325 = 66.5

after 325 minutes of use, plan B will cost more.